OpenStudy (anonymous):

how to solve this? Ahmad takes 6 days to complete his work, ali takes 3 days. how much time will be taken if both work together?

OpenStudy (anonymous):

2 days As Ahmad would have done 2/6 of his work and Ali would have done 2/3. 2/3 + 2/6 = 6/6.

OpenStudy (wolf1728):

There is a formula Total = (Time 1 • Time 2) / (Time 1 + Time 2) Total = (6 • 3) / (6 + 3) Total = 18 / 9 Total = 2 days

OpenStudy (anonymous):

thanx